# Beyond Incompatibility: Trade-offs between Mutually Exclusive Fairness Criteria in Machine Learning and Law

### v5

@@ Abstract
Fair and trustworthy AI is becoming ever more important in both machine learning and legal domains. One important consequence is that decision makers must seek to guarantee a ‘fair’, i.e., non-discriminatory, algorithmic decision procedure. However, there are several competing notions of algorithmic fairness that have been shown to be mutually incompatible under realistic factual assumptions. This concerns, for example, the widely used fairness measures of ‘calibration within groups’ and ‘balance for the positive/negative class’. In this paper, we present a novel algorithm (FAir Interpolation Method: FAIM) for continuously interpolating between these three fairness criteria. Thus, an initially unfair prediction can be remedied to, at least partially, meet a desired, weighted combination of the respective fairness conditions. We demonstrate the effectiveness of our algorithm when applied to synthetic data, the COMPAS data set, and a new, real-world data set from the e-commerce sector. Finally, we discuss to what extent FAIM can be harnessed to comply with conflicting legal obligations. The analysis suggests that it may operationalize duties in traditional legal fields, such as credit scoring and criminal justice proceedings, but also for the latest AI regulations put forth in the EU, like the Digital Markets Act and the recently enacted AI Act.
@@
{highlights}
Interpolation between mutually exclusive fairness criteria via optimal transport
Presentation of a novel method (FAIM) to reach any solution on the fairness simplex
Discussion on how FAIM can be harnessed to comply with conflicting legal obligations
@@ 1 Introduction
Fairness and non-discrimination have increasingly moved to the center of AI research and policy in recent years [77, 32, 114, 86, 16, 116]. While studies on bias and discrimination perpetuated and exacerbated by computer systems have a long pedigree [36, 54], recent advances in the AI space, including generative models, have brought an increased sense of urgency to this field. As machine learning (ML) progressively permeates all sectors of society, responsible and fair AI development is more crucial than ever [25, 44, 99]. This is particularly important in sensitive applications such as healthcare, hiring, credit, and law enforcement, where biased AI can have significant, real-world consequences.
@@
Addressing fairness and non-discrimination has become a key concern for the AI community, not only but also for instrumental reasons: it enhances the robustness and generalizability of AI models. Biased models are likely to perform poorly when faced with data that differs from their training set, limiting their applicability and effectiveness. By prioritizing diversity and fairness, researchers can create AI systems that are more accurate and reliable across a variety of contexts and populations. This broad applicability seems essential for the long-term success and acceptance of AI technologies.
@@
The devil, however, is in the details. For instance, in a much-noted incident, Google’s Gemini generated pictures of black and racially diverse US founding fathers, popes, and Nazi soldiers – in striking contrast to historical fact [60]. This example highlights that it is often difficult, and sometimes even impossible, to enforce a unidimensional fairness metric for appropriate model behavior. Rather, in many settings, different varieties of fairness have to be traded off against one another, and against the desire for (historical) accuracy, performance, and other competing normative goals.
@@
These inherent complexities and trade-offs will only become more important as specific legal provisions increasingly mandate fairness and nondiscrimination in AI [74, 51, 108]. In this context, discrimination against protected groups in decisions facilitated by ML models remains a key challenge. Non-discrimination legislation applying to biased AI systems does exist in the US and the EU, but loopholes and enforcement problems remain [6, 45, 115, 106, 104]. The new EU AI Act [^1] The final version of can be found under https://www.europarl.europa.eu/doceo/document/TA-9-2024-0138-FNL-COR01_EN.pdf compels AI developers to comprehensively assess and mitigate risks to fundamental rights, such as non-discrimination, stemming from high-risk AI systems and foundation models (Art. 9, 55 AI Act) [101, 29, 46]. This is backed up by a new framework for AI liability adopted by the EU legislator [47]. Moreover, the recently enacted EU Digital Markets Act (DMA [^2] Regulation (EU) 2022/1925 of the European Parliament and of the Council of 14 September 2022 on contestable and fair markets in the digital sector, OJ 2022, L265/1.) now compels large online platforms to engage in ‘fair and non-discriminatory’ ranking (Article 6(5) DMA), with a significant impact on the AI systems used for these tasks, which are key to the digital economy.
@@
Our paper develops one such algorithm allowing decision makers to interpolate between three different incompatible fairness criteria singled out for their importance in previous research. These three fairness criteria are described below. There are various ways in which unfairness, i.e., discrimination against certain groups  [36], may enter an ML model [15]. For example, the model may perpetuate historical biases present in the training data; use non-representative training data; or optimize for a target variable which unwittingly encodes differential access of individuals or groups [6]. The result of such biased modeling may be that false positive or false negative rates differ between groups of individuals. Specific fairness metrics such as equalized odds, quantifying the difference of false positive and false negative rates between groups  [53], are used to describe how much a classifier systematically rewards or punishes one group more than another with respect to ground truth. Even if a classifier is fair in terms of equalized odds, its prediction scores can lead to unfair outcomes if they are uncalibrated, as decision makers are not able to reliably interpret its output [109, 62]. [^3] A classifier is calibrated when, given a score $s$, $s$-fraction of all individuals receiving score $s$ are ground truth positive [22]. What is worse, if a classifier emits scores that are differently calibrated between groups, identical scores have different meanings for different protected groups, again leading to unfair outcomes [67].
@@
Given that equality of false positive rates between groups, equality of false negative rates between groups, and calibration between groups are all important, socially and even legally desirable fairness criteria [21], it would be ideal if a classifier could satisfy all three. Unfortunately, previous work shows that simultaneously fulfilling all three of these fairness criteria is impossible, except under highly constrained circumstances such as when a classifier achieves 100% accuracy with each instance receiving a prediction score of exactly 0 or 1 [20, 67, 87]. Pleiss et al. [87] attempt to find a relaxed notion of error rate difference that can be simultaneously achieved alongside group calibration. However, they demonstrate that one may only simultaneously achieve group calibration with a relaxed version of one of either equalized false positive or false negative rates. Furthermore, they show that classifiers simultaneously satisfying two of these three fairness criteria must engage in the equivalent of randomizing a certain percentage of predictions. Thus, the authors suggest that practitioners should only attempt to satisfy one of the three fairness criteria based on importance, instead of trying to simultaneously fulfill a relaxed notion of two of them. Often, however, all three fairness notions embody valid concerns of different groups [21]. Merely satisfying one of them may not do justice to the various interests and rights of the groups, and, as mentioned above, might even amount to unjustified discrimination if the law demands a balance rather than an absolute priority of one criterion.
@@
Building on the assumption that decision makers should determine, within the constraints of the law, which combination of the three fairness criteria is most crucial towards a fair result in practice, our work introduces a new post-processing algorithm called FAir Interpolation Method (FAIM). It allows decision makers to consciously and continuously interpolate between fairness criteria. The algorithm is based on the notion that, given a classifier, a representative evaluation data set, and the corresponding prediction scores for each group, any one of the three fairness criteria could be satisfied by applying a certain fairness criteria specific score mapping function mapping scores from the distribution outputted by the model to a fair score distribution. Given three such fairness criteria specific score distributions, FAIM uses optimal transport to continuously interpolate between them. This leads to a mapping function that can be applied to new classifier predictions, post hoc, to achieve a desired, weighted combination of the respective fairness criteria. FAIM cannot be used to fully satisfy all three fairness criteria simultaneously, but it does give the practitioner a powerful tool to make a continuous compromise along the fairness criteria simplex, and thus to account for different concerns, and legal obligations, each fairness criterion encompasses.
@@
The remainder of the article is organized as follows. First, related work is outlined in detail. Next, the mathematical foundations regarding fairness criteria incompatibility and optimal transport are addressed. Based on this, FAIM is introduced, and tested empirically in three experiments. In the first experiment, FAIM is applied to a synthetic classifier that produces two distinct score distributions for two respective groups. Here the properties of FAIM are elucidated under controlled conditions. In the second experiment, FAIM is tested on the COMPAS data set, on which the COMPAS algorithm famously scores individuals for recidivism risk. In the last experiment, FAIM is applied to product rankings from the European e-commerce platform Zalando.444www.zalando.de Here, we evaluate our model in terms of typical fairness problems arising in e-commerce rankings. The discussion takes up the results of the experiments and adds a specific legal perspective. We show how FAIM may be harnessed to trade off conflicting legal requirements across a wide variety of domains: credit scoring, criminal justice decisions, and fair rankings according to the DMA and the AI Act.
@@ 2 Related Work
The use of optimal transport methods in algorithmic fairness has become increasingly popular  [27, 31, 41, 112, 19, 18], applying optimal transport to trade-offs between individual vs group fairness, and accuracy vs fairness. Optimal transport is a mathematical tool that provides a very natural notion of distance of two probability distributions (the Wasserstein or earth mover distance) and, moreover, gives an optimal (in a specific sense) way to translate between these distributions by means of an optimal transport map or at least an optimal transport plan, the latter requiring some form of randomization. In [112], particularly, an important role is played by the Wasserstein-2 barycenter of several probability distributions and the displacement interpolation that allows to continuously interpolate between various distributions. We make use of these techniques in the present article.
@@
The literature at the intersection of law and computer science has, in the past few years, increasingly discussed how legal non-discrimination requirements can be integrated into code, and vice versa  [112, 45, 105, 106, 115, 37, 103, 104, 6, 90, 56, 80]. This strand of research has focused particularly on remedying discrimination in criminal justice proceedings [8, 98, 76, 93, 64, 28, 59, 89] and in other high-stakes AI decisions [102], such as employment contexts [90], credit scoring [10], or university admissions [112]. More specifically, Wachter et al. [105] rightly point out that the reliance on fairness metrics based on ground truth may perpetuate biases if that ground truth itself is skewed against protected groups. This is a concern, for example, in criminal justice settings [8, 5]. We incorporate this constraint into the application perimeter of our algorithm (see Part 6.3). Conversely, in the e-commerce sector, numerous publications deal with the emerging competition law framework for digital markets, for example the DMA [30, 48, 13, 14, 73, 88]. However, they do not, to our knowledge, specifically develop algorithms to solve legal questions arising in this field. The same holds true for the AI Act, where the literature, so far, focuses almost exclusively on the legal, economic and policy issues [101, 74, 50, 55, 81, 92] or on explainability questions [83, 97].
@@ 3.1 Incompatibility of Fairness Criteria
For the purpose of deriving the incompatibility theorem, it need not be assumed that the ground-truth dependent quantities $n_{t},x_{t},y_{t}$ be known; rather, whatever their values might be, the fairness criteria proposed in the next paragraph are never met except in case of perfect prediction or equal base rates. Here, by base rate for group $t$ we mean the ratio $n_{t}/N_{t}$, and perfect prediction means that the score of each individual is 1 when she is positive and 0 when she is negative.
@@
    $x_{1}=x_{2}=:x.$
@@
Let us illustrate the three requirements for the case of recidivism prediction. The members of each group (say, black and white offenders) are given a score that is intended to reflect the probability of the respective individual to recidivate within a given time period. Criterion A) then requires the following: The average score within group $t$ should equal the ratio of actual recidivists in that group; so if, say, half of the members of the white group would turn out recidivist, then the average score given to white individuals should be $0.5$. Concerning B), given a non-recidivist individual, it is desirable that their score be independent of their group membership. It would have to be considered highly unfair if a black person received a higher score, and thus faced a higher likelihood of detention, than a white person, given that both would actually not commit an offense if released. This, in fact, is at the heart of the controversy around the COMPAS algorithm. Note that B) only considers scores on an average level: The average score among all truly non-recidivist individuals should be equal between groups. Of course, this tells nothing about higher order statistical moments, such as the variance of the respective score distributions. The discussion for criterion C) is analogous.
@@
which is zero if and only if either (1) $n_{1}\neq 0\neq n_{2}$and $\frac{N_{1}-n_{1}}{n_{1}}=\frac{N_{2}-n_{2}}{n_{2}}$, which is the case of equal base rates; or (2) one of $n_{1}$ or $n_{2}$ is zero, in which case (assuming $N_{1},N_{2}>0$) also the other one is zero, i.e., there are no positive individuals at all. Here, the choice $x=0$ and $y$ arbitrary yields a fair assignment according to Eq. 1. If there are no positive individuals, then everyone should get score zero, which is thus an instance of perfect prediction. For the first case, there are infinitely many solutions, one of which is $x=y=\frac{n_{1}}{N_{1}}=\frac{n_{2}}{N_{2}}$ (as mentioned in Kleinberg et al. [67]). This is realizable by assigning the same score to everyone. (A mathematically simpler solution would be $x=0,y=1$, which however requires the algorithm to know who is positive and who isn’t and thus again requires perfect prediction.)
@@
In summary, we have thus verified Theorem 1.1, the main result, of Kleinberg et al. [67], which we paraphrase as follows:

### v6

@@ Abstract
Fair and trustworthy AI is becoming ever more important in both machine learning and legal domains. One important consequence is that decision makers must seek to guarantee a ‘fair’, i.e., non-discriminatory, algorithmic decision procedure. However, there are several competing notions of algorithmic fairness that have been shown to be mutually incompatible under realistic factual assumptions. This concerns, for example, the widely used fairness measures of ‘calibration within groups’ and ‘balance for the positive/negative class,’ which relate to accuracy, false negative and false positive rates, respectively. In this paper, we present a novel algorithm (FAir Interpolation Method: FAIM) for continuously interpolating between these three fairness criteria. Thus, an initially unfair prediction can be remedied to meet, at least partially, a desired, weighted combination of the respective fairness conditions. We demonstrate the effectiveness of our algorithm when applied to synthetic data, the COMPAS data set, and a new, real-world data set from the e-commerce sector. We provide guidance on using our algorithm in different high-stakes contexts, and we discuss to what extent FAIM can be harnessed to comply with conflicting legal obligations. The analysis suggests that it may operationalize duties in traditional legal fields, such as credit scoring and criminal justice proceedings, but also for the latest AI regulations put forth in the EU, like the Digital Markets Act and the recently enacted AI Act.
@@ 1 Introduction
Fairness and non-discrimination have increasingly moved to the center of AI research and policy in recent years [84, 34, 122, 94, 18, 124]. While studies on bias and discrimination perpetuated and exacerbated by computer systems have a long pedigree [38, 59], recent advances in the AI space, including generative models, have brought an increased sense of urgency to this field [10, 75, 56, 55]. As machine learning (ML) progressively permeates all sectors of society, responsible and fair AI development is more crucial than ever [27, 48, 107]. This is particularly important in sensitive applications such as healthcare, hiring, credit, and law enforcement, where biased AI can have significant, real-world consequences.
@@
Addressing fairness and non-discrimination has become a key concern for the AI community, not only but also for instrumental reasons: while furthering social justice, they also enhance the robustness and generalizability of AI models. Biased models are likely to perform poorly when applied to certain subgroups or when faced with data that differs from their training set, limiting their applicability and effectiveness [62, 2, 46]. By prioritizing diversity and fairness, researchers can create AI systems that are more accurate and reliable across a variety of contexts and populations. This broad applicability seems essential for the long-term success and acceptance of AI technologies.
@@
The devil, however, is in the details. For instance, in a much-noted incident, Google’s Gemini model generated pictures of black and racially diverse US founding fathers, popes, and Nazi soldiers – in striking contrast to historical fact [66]. This example highlights that it is often difficult, and sometimes even impossible, to enforce a unidimensional fairness metric for appropriate model behavior. Rather, in many settings, different varieties of fairness have to be traded off against one another, and against the desire for (historical) accuracy, performance, and other competing normative goals.
@@
These inherent complexities and trade-offs will only become more important as specific legal provisions increasingly mandate fairness and non-discrimination in AI [81, 55, 116]. In this context, discrimination against protected groups in decisions facilitated by ML models remains a key challenge. Non-discrimination legislation applying to biased AI systems does exist in the US and the EU, but loopholes and enforcement problems remain [7, 49, 123, 114, 112]. The new EU AI Act [^1] Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence (Artificial Intelligence Act), OJ L, 2024/1689. compels AI developers to comprehensively assess and mitigate risks to fundamental rights, such as non-discrimination, stemming from high-risk AI systems and large general-purpose models (Art. 9, 55 AI Act) [109, 31, 50]. This is backed up by a new framework for AI liability adopted by the EU legislator [51]. Moreover, the EU Digital Markets Act (DMA [^2] Regulation (EU) 2022/1925 of the European Parliament and of the Council of 14 September 2022 on contestable and fair markets in the digital sector, OJ 2022, L265/1.) now compels large online platforms to engage in ‘fair and non-discriminatory’ ranking (Article 6(5) DMA), with a significant impact on the AI systems used for these tasks, which are key to the digital economy.
@@
Our paper develops one such algorithm allowing decision makers to interpolate between three different incompatible fairness criteria singled out for their importance in previous research. These three fairness criteria are described below. There are various ways in which unfairness, i.e., discrimination against certain groups  [38], may enter an ML model [17]. For example, the model may perpetuate historical biases present in the training data; use non-representative training data; or optimize for a target variable which unwittingly encodes differential access of individuals or groups [7]. The result of such biased modeling may be that false positive or false negative rates differ between groups of individuals. Specific fairness metrics such as equalized odds, quantifying the difference of false positive and false negative rates between groups  [58], are used to describe how much a classifier systematically rewards or punishes one group more than another with respect to ground truth. Even if a classifier is fair in terms of equalized odds, its prediction scores can lead to unfair outcomes, however, if they are uncalibrated, as decision makers are not able to reliably interpret its output [117, 68]. [^3] A classifier is calibrated when, given a score $s$, $s$-fraction of all individuals receiving score $s$ are ground truth positive [24]. What is worse, if a classifier emits scores that are differently calibrated between groups, identical scores have different meanings for different protected groups, again leading to unfair outcomes [73].
@@
Given that equality of false positive rates between groups, equality of false negative rates between groups, and calibration between groups are all important, socially and even legally desirable fairness criteria [23], it would be ideal if a classifier could satisfy all three. Unfortunately, previous work shows that simultaneously fulfilling all three of these fairness criteria is impossible, except under highly constrained circumstances such as when a classifier achieves 100% accuracy with each instance receiving a prediction score of exactly 0 or 1 [22, 73, 95]. Merely focusing on one of these metrics may not do justice to the various interests and rights of the groups, and, as mentioned above, might even amount to unjustified discrimination if the law demands a balance rather than an absolute priority of one criterion.
@@
Building on the assumption that decision makers should determine, within the constraints of the law, which combination of the three fairness criteria is most crucial towards a fair result in practice, our work introduces a new post-processing algorithm called FAir Interpolation Method (FAIM). It allows decision makers to consciously and continuously interpolate between fairness criteria. The algorithm is based on the notion that, given a classifier, a representative evaluation data set, and the corresponding prediction scores for each group, any one of the three fairness criteria could be (partially) satisfied by applying a certain fairness-criteria-specific score mapping function, which maps scores from the distribution outputted by the model to a fair score distribution. Given three such fairness-criteria-specific score distributions, FAIM uses optimal transport to continuously interpolate between them. This leads to a mapping function that can be applied to new classifier predictions, post hoc, to achieve a desired, weighted combination of the respective fairness criteria. FAIM cannot be used to fully satisfy all three fairness criteria simultaneously, but it does give the practitioner a powerful tool to make a continuous compromise along the fairness criteria simplex, and thus to account for different concerns, and legal obligations, each fairness criterion encompasses.
@@
The remainder of the article is organized as follows. First, related work is outlined in detail. Next, the mathematical foundations regarding fairness criteria incompatibility and optimal transport are addressed. Based on this, FAIM is introduced, and tested empirically in three experiments. In the first experiment, FAIM is applied to a synthetic classifier that produces two distinct score distributions for two respective groups. Here the properties of FAIM are elucidated under controlled conditions. In the second experiment, FAIM is tested on the COMPAS data set, on which the COMPAS algorithm famously scores individuals for recidivism risk. In the last experiment, FAIM is applied to product rankings from the European e-commerce platform Zalando.444www.zalando.de Here, we evaluate our model in terms of typical fairness problems arising in e-commerce rankings. The discussion takes up the results of the experiments and adds a specific legal perspective. We show how FAIM may be harnessed to trade off conflicting legal requirements across a wide variety of domains: credit scoring, criminal justice decisions, and fair rankings according to the DMA and the AI Act. The paper also provides guidance on how the different fairness criteria can be weighted and traded off against one another in a variety of high-stakes settings, ranging from recidivism prediction, credit scoring and e-commerce to hiring, college admissions, and healthcare.
@@ 2 Related Work
While optimal transport provides a mathematically principled framework for interpolating between fairness criteria, it remains important to demonstrate why this approach was selected and how it compares to potential alternatives. In the current literature, there are few established methods for continuously navigating between different fairness notions. Most existing approaches treat fairness criteria as binary constraints that must be either fully satisfied or not, without allowing for meaningful intermediate solutions. The optimal transport framework, in contrast, provides a mathematically rigorous way to quantify and optimize the trade-offs between competing fairness criteria. By leveraging the theory of optimal transport, our approach guarantees that the interpolation between fairness criteria is optimal in terms of minimizing the cost of transforming one fairness regime into another, while preserving the underlying probabilistic structure of the problem[120].
Some previous work has looked into the possibility of simultaneously fulfilling several fairness constraints in modified ways. Pleiss et al. [95] attempt to find a relaxed notion of error rate difference that can be simultaneously achieved alongside group calibration. However, they demonstrate that one may only simultaneously achieve group calibration with a relaxed version of one of either equalized false positive or false negative rates. Furthermore, they show that classifiers simultaneously satisfying two of these three fairness criteria must engage in the equivalent of randomizing a certain percentage of predictions. Thus, the authors suggest that practitioners should only attempt to satisfy one of the three fairness criteria based on importance, instead of trying to simultaneously fulfill a relaxed notion of two of them. Often, however, all three fairness notions embody valid concerns of different groups [23]. This is what FAIM acknowledges.
One might consider alternative approaches such as quota-based systems, which have historically been used in domains like college admissions or hiring to balance competing fairness objectives. However, such methods suffer from several limitations. First, quota systems are typically applicable only to specific scenarios involving discrete allocation decisions, while our FAIM framework can be applied broadly across different types of algorithmic decision-making systems. Second, quota-based approaches often implement rather crude trade-offs, whereas our optimal transport method allows for more nuanced, continuous adjustments between different fairness criteria. Additionally, quota systems may face legal challenges in many jurisdictions, particularly under anti-discrimination laws in the US and the European Union, which generally prohibit rigid quota systems in favor of more flexible approaches to ensuring fairness [49, 39]. A theoretical alternative might involve direct linear interpolation between different fairness metrics. However, such an approach would fail to account for the geometric structure of the probability spaces involved and could lead to implementations that violate basic probabilistic constraints. Our optimal transport-based framework, by contrast, explicitly preserves the probabilistic nature of the problem while providing a mathematically principled way to navigate the space of fairness criteria. This ensures that intermediate solutions remain well-defined probability distributions and maintain desirable statistical properties throughout the interpolation process.
@@
More generally, the use of optimal transport methods in algorithmic fairness has become increasingly popular  [29, 33, 44, 120, 21, 20], applying optimal transport to trade-offs between individual vs group fairness, and accuracy vs fairness. Optimal transport is a mathematical tool that provides a very natural notion of distance of two probability distributions (the Wasserstein or earth mover distance) and, moreover, gives an optimal (in a specific sense) way to translate between these distributions by means of an optimal transport map or at least an optimal transport plan, the latter requiring some form of randomization. In [120], particularly, an important role is played by the Wasserstein-2 barycenter of several probability distributions and the displacement interpolation that allows to continuously interpolate between various distributions. We make use of these techniques in the present article.
@@
The literature at the intersection of law and computer science has, in the past few years, increasingly discussed how legal non-discrimination requirements can be integrated into code, and vice versa  [120, 49, 113, 114, 123, 40, 111, 112, 7, 98, 61, 88]. This strand of research has focused particularly on remedying discrimination in criminal justice proceedings [9, 106, 83, 101, 70, 30, 65, 97] and in other high-stakes AI decisions [110], such as employment contexts [98], credit scoring [12], or university admissions [120]. More specifically, Wachter et al. [113] rightly point out that the reliance on fairness metrics based on ground truth may perpetuate biases if that ground truth itself is skewed against protected groups. This is a concern, for example, in criminal justice settings [9, 6]. We incorporate this constraint into the application perimeter of our algorithm (see Part 6.4). Conversely, in the e-commerce sector, numerous publications deal with the emerging competition law framework for digital markets, for example the DMA [32, 52, 15, 16, 80, 96]. However, they do not, to our knowledge, specifically develop algorithms to solve legal questions arising in this field. The same holds true for the AI Act, where the literature, so far, focuses almost exclusively on the legal, economic and policy issues [109, 81, 54, 60, 89, 100] or on explainability questions [91, 105].
@@ 3.1 Incompatibility of Fairness Criteria
For the purpose of deriving the incompatibility theorem, it need not be assumed that the ground-truth dependent quantities $n_{t},x_{t},y_{t}$ be known; rather, whatever their values might be, the fairness criteria proposed in the next paragraph are never simultaneously met except in case of perfect prediction or equal base rates. Here, by base rate for group $t$ we mean the ratio $n_{t}/N_{t}$, and perfect prediction means that the score of each individual is 1 when she is positive and 0 when she is negative.
@@
    $x_{1}=x_{2}=:x,$
    where the symbol ‘$=$:’ denotes ‘equal by definition’, i.e., $x$ is defined as the common value of $x_{1}$ and $x_{2}$ in the balanced case.
@@
Let us illustrate the three requirements for the case of recidivism prediction. The members of each group (say, black and white offenders) are given a score that is intended to reflect the probability of the respective individual to recidivate within a given time period. Criterion A) then requires the following: The average score within group $t$ should equal the ratio of actual recidivists in that group; so if, say, half of the members of the white group would turn out recidivist, then the average score given to white individuals should be $0.5$. Concerning B), given a non-recidivist individual, it is desirable that their score be independent of their group membership. It would have to be considered highly unfair if a black person received a higher score, and thus faced a higher likelihood of detention, than a white person, given that both would actually not commit an offense if released. This, in fact, is at the heart of the controversy around the COMPAS algorithm. Note that B) only considers scores on an average level: The average score among all truly non-recidivist individuals should be equal between groups. Of course, this tells nothing about higher order statistical moments, such as the standard deviation of the respective score distributions. The discussion for criterion C) is analogous.
@@
which is zero if and only if either (1) $n_{1}$and $n_{2}$ are both not zeroand $\frac{N_{1}-n_{1}}{n_{1}}=\frac{N_{2}-n_{2}}{n_{2}}$, which is the case of equal base rates; or (2) one of $n_{1}$ or $n_{2}$ is zero, in which case (assuming $N_{1},N_{2}>0$) also the other one is zero, i.e., there are no positive individuals at all. Here, the choice $x=0$ and $y$ arbitrary yields a fair assignment according to Eq. 1. If there are no positive individuals, then everyone should get score zero, which is thus an instance of perfect prediction. For the first case, there are infinitely many solutions, one of which is $x=y=\frac{n_{1}}{N_{1}}=\frac{n_{2}}{N_{2}}$ (as mentioned in Kleinberg et al. [73]). This is realizable by assigning the same score to everyone. (A mathematically simpler solution would be $x=0,y=1$, which however requires the algorithm to know who is positive and who isn’t and thus again requires perfect prediction.)
@@
In summary, we have thus verified Theorem 1.1, the main result of Kleinberg et al. [73], which we paraphrase as follows: